# LongFaith: Enhancing Long-Context Reasoning in LLMs with Faithful Synthetic Data

### v1

@@ Abstract
Cehao Yang1,2111Equal contribution., Xueyuan Lin1,2111Equal contribution., Chengjin Xu1111Equal contribution., Xuhui Jiang1, Shengjie Ma1, Aofan Liu1, Hui Xiong2222Corresponding authors., Jian Guo1222Corresponding authors., 1IDEA Research, International Digital Economy Academy 2Artificial Intelligence Thrust, Hong Kong University of Science and Technology (Guangzhou) {yangcehao,linxueyuan,xuchengjin,jiangxuhui,mashengjie,liuaofan,guojian}@idea.edu.cn {cyang289,xlin058}@connect.hkust-gz.edu.cn, xionghui@ust.hk
@@ 1 Introduction
Long-context processing ability has emerged as a significant challenge for large language models (LLMs) Shi et al. (2023); Liu et al. (2024); Wu et al. (2024); Levy et al. (2024), especially arises when models process extensive textual information, making it hard to recognize relevant evidence and address downstream tasks such as question answering (QA), summarization, and complex reasoning Bai et al. (2023, 2024b); Zhang et al. (2024d); Hsieh et al. (2024); Yen et al. (2024). A variety of model-centric methods have been proposed to extend the length of context windows in LLMs Chen et al. (2023a, b); Peng et al. (2023); Han et al. (2024); Ding et al. (2024). Additionally, many data-centric methods, such as synthesizing long-context understanding instructions for fine-tuning, have gained attention for enhancing models’ ability to handle and utilize extended contexts Xiong et al. (2023); An et al. (2024); Fu et al. (2024); Bai et al. (2024a); Li et al. (2024b); Gao et al. (2024); Chen et al. (2024); Zhang et al. (2024c); Li et al. (2024a).

### v2

@@ Abstract
Cehao Yang1,2111Equal contribution., Xueyuan Lin1,2,4111Equal contribution., Chengjin Xu1,3111Equal contribution., Xuhui Jiang1,3, Shengjie Ma1, Aofan Liu1, Hui Xiong2222Corresponding authors., Jian Guo1222Corresponding authors. 1IDEA Research, International Digital Economy Academy 2Hong Kong University of Science and Technology (Guangzhou) 3DataArc Tech Ltd. 4Hithink RoyalFlush Information Network Co., Ltd {yangcehao,linxueyuan,xuchengjin,jiangxuhui,mashengjie,liuaofan,guojian}@idea.edu.cn {cyang289,xlin058}@connect.hkust-gz.edu.cn, xionghui@ust.hk
@@ 1 Introduction
Long-context processing ability has emerged as a significant challenge for large language models (LLMs) Shi et al. (2023); Liu et al. (2024); Wu et al. (2024); Levy et al. (2024), especially arises when models process extensive textual information, making it hard to recognize relevant evidence and address downstream tasks such as question answering (QA), summarization, and complex reasoning Bai et al. (2023, 2024b); Zhang et al. (2024d); Hsieh et al. (2024); Yen et al. (2024). A variety of model-centric methods have been proposed to extend the length of context windows in LLMs Chen et al. (2023a, b); Peng et al. (2023); Han et al. (2024); Ding et al. (2024). Additionally, many data-centric methods, such as synthesizing long-context understanding instructions for fine-tuning, have gained attention for enhancing models’ ability to handle and utilize extended contexts Xiong et al. (2023); An et al. (2024); Fu et al. (2024); Bai et al. (2024a); Li et al. (2024b); Gao et al. (2024); Chen et al. (2024); Zhang et al. (2024c); Li et al. (2024a); Jiang et al. (2025).
@@ 4.5 Analysis
Llama-3.1-8B LongBench(S) LongBench(M)
EM F1 EM F1
   w/ 10 docs 11.2 31.7 39.7 52.2
   w/ 20 docs 15.4 35.3 44.8 55.6
   w/ 30 docs 16.0 36.0 46.1 56.7
   w/ 40 docs 16.9 37.8 47.2 58.1
Table 5: Effectiveness of LongFaith on LongBench as context length increases.
Effectiveness of LongFaith as Context Length Increases.
We further investigate the impact of increasing context length on the performance of models trained with the LongFaith pipeline. Specifically, we manipulate the number of documents included in the training set MuSiQue to simulate varying context lengths during training. By incrementally increasing the number of documents from 10 to 40, we assess how the model’s reasoning ability scales when exposed to longer input contexts. The results, as reported in Tab. 5, show a consistent improvement in both EM and F1 scores across both LongBench(S) and LongBench(M) as the context length increases. This indicates that LongFaith effectively leverages additional contextual information, enhancing the model’s comprehension and reasoning capabilities. These findings validate the scalability of the LongFaith framework in handling long-context scenarios, highlighting its potential for applications that require deep reasoning over extensive inputs.